Daily rate is $39 + 0.20 per mile, Rent car for 1 day, the bill is $95 before taxes. How many miles did you drive.
Am I setting this equation up right? Here is what I have.
39+ 0.20 (1) = 95 If not how??

Answer by mananth(16949)

$39 + 0.20 per mile,
if you drive only 1 mile it will cost 39.20
if you drive x miles it will cost (%
$39 + 0.20(x)=95
0.2x=95-39
0.2x=56
/0.2
x=56/0.2
x=280 miles